A sample that contains 3.98 mol of a gas at
255 K has a pressure of 0.621 atm. What is
the volume?
1
Expert's answer
2020-04-28T12:19:35-0400

"P=0.621\ atm,V=?,T=255K,n=3.98"

Using ideal gas equation,"PV=nRT" and "R=0.0821\ L\ atm\ mol^{-1}K^{-1}"

On substituting these values,

"0.621\times V=0.0821\times 3.98\times 255\implies V=134\ L"